### v1

# Rehearsal-free Federated Domain-incremental Learning

Abstract
We introduce a rehearsal-free federated domain incremental learning framework, RefFiL, based on a global prompt-sharing paradigm to alleviate catastrophic forgetting challenges in federated domain-incremental learning, where unseen domains are continually learned. Typical methods for mitigating forgetting, such as the use of additional datasets and the retention of private data from earlier tasks, are not viable in federated learning (FL) due to devices’ limited resources. Our method, RefFiL, addresses this by learning domain-invariant knowledge and incorporating various domain-specific prompts from the domains represented by different FL participants. A key feature of RefFiL is the generation of local fine-grained prompts by our domain adaptive prompt generator, which effectively learns from local domain knowledge while maintaining distinctive boundaries on a global scale. We also introduce a domain-specific prompt contrastive learning loss that differentiates between locally generated prompts and those from other domains, enhancing RefFiL’s precision and effectiveness. Compared to existing methods, RefFiL significantly alleviates catastrophic forgetting without requiring extra memory space, making it ideal for privacy-sensitive and resource-constrained devices.
1 Introduction
Federated Learning (FL) is a remarkable approach that enables various participants, also referred to as clients, to collaboratively train a global machine learning model while keeping their data decentralized, maintaining privacy and leveraging diverse data sources [24, 9]. A significant limitation of current FL research lies in its primary focus on scenarios with static data distributions [20, 12, 17, 13, 34], which remain unchanged throughout the training process [8, 7]. This limitation is particularly pronounced as it does not reflect the dynamic nature of real-world data. Recent studies have therefore shifted towards Federated Continual Learning (FCL), especially in class-incremental data scenarios [8, 7, 38, 23, 2], which struggles significantly with the critical problem of catastrophic forgetting [30], where model rapidly forgets previously learned knowledge upon encountering new data. In this paper, we explore more challenging domain-incremental data scenarios.
Figure 1: Key steps of RefFil framework. Panel (a) shows the common setting in existing FCL works, characterized by a cliff-style data transition. Panel (b) shows our approach, where a subset of participants gradually transitions to new tasks. Panel (c) provides an overview example of a key step in our methodology: the 1st participant processes new domain data using global prompts from the 2nd to $m$-th participants and local prompts, enhancing robustness by aligning the model’s predictions across diverse domain prompts as inputs.
Various methods have been proposed to address catastrophic forgetting. Rehearsal-based approaches, such as those described in [30], mitigate forgetting by replaying a selection of old task data during new task training. Alternatively, regularization-based methods, like Knowledge Distillation (KD) [21, 30, 3, 8], guide the model’s optimization using historical data or models. Network expansion techniques [18, 11, 29] adapt the model structure to incorporate new features but their practicality in memory-constrained FL environments is limited. Recent research like L2P [37] and DualPrompt [36] explores prompt-based, rehearsal-free domain-incremental learning. However, these methods still do not fully address the challenge of using global domain information to enhance local model robustness in Federated Domain-Incremental Learning (FDIL).
Strategies for mitigating catastrophic forgetting in domain-incremental tasks are distinct from those employed in class-incremental scenarios, where the substantial differences between classes make it easier for models to distinguish decision boundaries. In domain-incremental contexts, however, the smaller feature differences between domains make it harder to distinguish decision boundaries. Thus, the priority shifts towards acquiring domain-invariant knowledge, an effort pivotal to enhancing model robustness. This focus becomes particularly vital when models are required to make predictions across a spectrum of diverse domains. To date, research in this area is notably sparse, especially concerning the effective global utilization of heterogeneous domain information from all clients at various stages of their transition. Such utilization is key to mitigating catastrophic forgetting by strengthening the local models’ proficiency in learning domain-invariant knowledge in FDIL and across the broader expanse of FCL.
In this paper, we introduce RefFil, a rehearsal-free federated domain-incremental learning framework that harnesses a wide spectrum of global domain information to augment cross-domain prediction robustness by specifically enhancing the acquisition of domain-invariant knowledge. We first recognize the inherent challenges posed by the high autonomy of participants in FL, which introduces significant uncertainty in data distribution and, consequently, a substantial domain gap. Thus, derivating robust prompts becomes critical in learning domain-invariant knowledge.
Hence, firstly, we define a practical FDIL problem. To address the challenge of significant domain gaps in datasets from various resource-sensitive participants, we propose a client-wise domain adaptive prompt generator, which creates personalized, instance-level prompts for each participant’s data, incorporating domain adaptation knowledge for global sharing. Then, to effectively break down the silos of domain-specific information in FL, we introduce a global prompt learning paradigm. This paradigm involves globally distributing and sharing clustered local prompts combined with local domain-invariant knowledge learning through prompting. By doing so, we aim to empower the local model to effectively differentiate between prompt features from various domains thereby enhancing its ability to learn domain-invariant knowledge from local data, which is crucial for boosting the model’s robustness in multi-domain predictions and significantly reducing its susceptibility to catastrophic forgetting. Finally, we introduce a novel domain-specific prompt contrastive learning method with temperature decay aimed at improving the local model’s ability to distinguish between semantically related and unrelated prompts. By seamlessly integrating these three key components, the RefFil framework emerges as a highly competitive method in the realm of FDIL.
2 Related Work
Most recently, several works have begun to concentrate on FCL, especially Federated Class-incremental Learning (FCIL), which involves sequentially learning a series of distinct, non-overlapping classes. The primary challenge in FCIL is catastrophic forgetting, where models tend to lose previously learned knowledge upon task transition. Approaches such as GFLC [8] and its extension LGA [7] use knowledge distillation, employing an old model to regularize the current model’s output and a loss function to mitigate global forgetting. Method TARGET [38] uses a global model to generate synthetic data with a global distribution for the current training task to reduce forgetting. CFed [23] addresses data unavailability by generating pseudo-labels on a surrogate dataset for knowledge distillation and proposes a server distillation mechanism to mitigate intra-task forgetting. These methods typically require storing old tasks or extra data, which is impractical in FL with resource-limited devices. MFCL [2] utilised GAN to generate data rather than save raw to achieve rehearsal-free learning from class-incremental data. Despite progress in the class-incremental scenario of FCL, there remains a notable gap in the domain-incremental setting, which focuses on learning domain-invariant knowledge to reduce forgetting, similar to domain adaptation [6], a different approach from FCIL.
To address the devices’ resource-constrained limitation in FDIL, the concept of prompt techniques, initially introduced in natural language processing (NLP) for task adaptation [22], offers a promising solution. These techniques append instructions to large pre-trained models, helping them use existing knowledge instead of learning from scratch. Applied to continual learning, approaches like L2P [37] and DualPrompt [36] use prompts to avoid the need for rehearsal. L2P uses a prompt pool for task-specific knowledge, replacing the rehearsal buffer. DualPrompt, designed for resource-limited devices, uses two prompt types: General-Prompt for general instructions and Expert-Prompt for task-specific guidance to replace maintaining a prompt pool. However, in FDIL, the challenge goes beyond learning local domain-invariant information; it involves using data from other participants to enhance model robustness against diverse domains. This paper aims to address this specific challenge, unexplored in current research.
To enhance local domain-invariant knowledge learning through the sharing of diverse domain information via global prompts in FDIL settings, we introduce RefFil, a rehearsal-free framework that leverages global prompts to share diverse domain information. As shown in Fig. 2, RefFil operates in task stages where participants engage in different domains. The process starts with a feature extractor generating a feature map from the input image, followed by our Client-wise Domain Adaptive Prompt (CDAP) generator. This generator encodes prompts using the feature map and a task ID. These encoded prompts, combined with the feature map, undergo classification involving computation of the cross-entropy loss and the Domain-specific Prompt Contrastive Learning (DPCL) loss with global prompts. Additionally, the feature map is merged with global prompts to compute the Global Prompt Learning (GPL) loss. All local prompts are sent to the central server to facilitate domain information sharing. For all preliminaries, including FDIL problem definition, please refer to Appendix A.
3 The Proposed Framework RefFil
Figure 2: Overview of the RefFil Framework. Each participant first encodes local prompts using the tokenized feature map and task ID embedding. These local prompts are then concatenated with the feature map to compute the loss $\mathcal{L}_{CE}$. Simultaneously, the feature map is combined with global prompts to calculate the loss $\mathcal{L}_{GPL}$, and the loss $\mathcal{L}_{DPCL}$ is determined between global and local prompts. Subsequently, all local prompts, along with the updated local models, are transmitted to the central server. The server then clusters the prompts domain-wise and aggregates the local models for distribution in the next training round.
Client-wise Domain Adaptive Prompt Generator. Most of the stages in FDIL training typically involve datasets from various domains, creating a substantial domain gap between participants’ local data. To navigate this challenge, generating refined prompts is essential for enabling participants to effectively distinguish and share prompts globally. Common approaches include maintaining a prompt pool, as seen in methods like L2P [37] and DualPrompt [36]. These strategies are similar to traditional continual learning methods, which rely on storing raw data or summarised representative information. However, a significant obstacle in FL is the limited memory resources of participant devices and privacy sensitivity. Additionally, as noted in the [16], a fixed-size prompt pool may not effectively encode domain-specific knowledge, and a scalable pool faces feasibility issues due to uncertainties in task requirements and computational constraints. This makes maintaining an extensive and adaptable prompt pool challenging for continuous learning in a federated context, where the unpredictability of tasks and resource limitations of devices must be considered.
Thus, inspired by the prompt encoder network structure proposed in [16], we propose a client-wise domain adaptive prompt generator. This generator is designed to create personalized prompts for each participant’s local data, embedding instance-level, fine-grained instructions with domain adaptation knowledge for global sharing. The CDAP generator $\mathcal{G}$ comprises several components: Layer Normalization (LN), a Multi-Layer Perceptron (MLP), a Cross-Client Domain Adaptation (CCDA) layer, and a Linear Transformation Layer (LT). This configuration enables the generation of adaptive local prompts $p_{m}$ for each participant $m$, which could be formulated as:
$P_{m}$ $=\operatorname{LT}\left(\operatorname{CCDA}\left(\operatorname{MLP}\left(\operatorname{LN}(\mathbf{I})^{\top}\right)\right);\phi(v)\right)^{\top}$ (1)
$=\left(\alpha_{v}\operatorname{CCDA}\left(\operatorname{MLP}\left(\operatorname{LN}(\mathbf{I})^{\top}\right)+\lambda_{v}\right)^{\top}\right)\in\mathbb{R}^{p\times d}\text{,}$
where $\mathbf{I}$ is a set of input tokens, $\phi$ represents a linear layer that predicts two types of affine transformation parameters, $\alpha$ and $\lambda$, from a conditional input embedding $v$. Initially, $v$ is normalized, transposed, from $(n+1)\times d$ to $d\times(n+1)$ and then passed through an MLP to generate $p$-dimensional outputs which create current task domain well-fitted and instance-level prompts tailored with domain-specific knowledge. To handle the data that vary significantly in distribution from various domains, we introduce a CCDA layer, enhancing the CDAP generator’s generalization and robustness via FL aggregation. Additionally, we incorporate a Feature-Wise Linear Modulation (LT) [28] framework to augment the generated prompts with additional instructional guidance for prediction. LT achieves this by applying an affine transformation to the instance-level prompts, using scaling $\alpha_{v}$ and shifting parameters $\lambda_{v}$, both conditioned on the embedding $v$ as $[\alpha_{v},\lambda_{v}]=\phi(v)$. For the conditional input embedding $v$, we designed a task-specific key embedding layer that is also employed to encode local task IDs, linking tasks with domain-specific data.
Global Prompt Sharing. To overcome domain information silos in FL and address the challenge of data quantity disparity among participants, we propose globally distributing local prompts in a balanced manner. This method prevents network overload in scenarios involving thousands of devices and mitigates biases due to uneven data distribution. Our solution involves averaging across all clients, ensuring equitable influence from each participant on the global model, regardless of their data volume. This strategy is crucial in situations where differences in data quantity between resource-rich and resource-poor participants could significantly skew the global model optimization direction. The formulation for the Local Prompts Group ($LPG$) is as follows:
$LPG_{m}=\left[\frac{1}{N}\sum_{j=1}^{N}P^{1}_{m,j},\dots,\frac{1}{N}\sum_{j=1}^{N}P^{k}_{m,j}\right]\in\mathbb{R}^{d}\text{,}$ (2)
then global prompts set $P^{g}$ could be summarised as:
$P^{g}=\left[LPG_{1},\dots,LPG_{m}\right]\text{.}$ (3)
However, generally, a significant portion of participants, 80% in our scenario, transition to training tasks in new domains, while a minority continues with old tasks. This leads to a domain-specific prompt imbalance, with most prompts coming from new domains. This imbalance may bias local model updates towards these majority domains.
Global Prompts Clustering. To tackle the prompt imbalance issue, directly averaging all prompts may lead to a loss of important domain-characterized features, as averaging might not fully capture the unique characteristics essential for effective learning in diverse domains. Therefore, we propose the use of global prompts clustering that filters and selects representative prompts for each class based on their semantic domain characteristics. We adopt FINCH [31], a parameter-free clustering method and particularly well-suited for large-scale prompts clustering with minimal computing resources, which is FL dynamic environment friendly. Additionally, its efficiency in clustering feature prototypes in FL has been previously demonstrated in [15]. FINCH operates on the principle that the nearest neighbor prompt of each sample can effectively support grouping. This is highly effective for distinguishing between prompts from different domains, as prompts from separate domains are unlikely to be nearest neighbors and are thus grouped separately. In contrast, similar prompts, likely from the same domain, are grouped together. Prompt similarity is evaluated using cosine similarity, where prompts are grouped by their smallest distance. This allows us to sort prompts into domain-specific sets efficiently. For each class $k$ of prompts, we then construct an adjacency matrix as:
$A^{k}(m,j)=\begin{cases}1,&\text{ if }j=c_{m}^{k}\text{ or }m=c_{j}^{k}\text{ or }c_{m}^{k}=c_{j}^{k}\\ 0,&\text{ otherwise }\end{cases}\text{,}$ (4)
where $c_{m}^{k}$ denotes the nearest neighbor (with the highest cosine similarity) of the class $k$ prompt from the $m$-th participant. Based on this clustering (Eq.4), we select representative prompts for each class from the embedding space. The final global clustered prompts are denoted as $\hat{P}^{g}$, are obtained as follows:
$\hat{P}^{g,k}=\left\{LPG_{m}\right\}_{m=1}^{C}\xrightarrow{\Psi}\left\{LPG_{m}\right\}_{m=1}^{N}\in\mathbb{R}^{N\times d}\text{,}$ (5)
$\hat{P}^{g}=\left[\hat{P}^{g,1},\dots,\hat{P}^{g,k}\right]\text{.}$
Finally, $C$ prompts are clustered by cluster model $\Psi$ into $N$ representatives of class $k$.
Personalised Prompt Generator Optimization. We expect to improve the local model’s ability to differentiate between semantically related and unrelated prompts, ensuring a clear class decision boundary by enhancing sensitivity to domain-specific nuances in prompts. Specifically, we encourage the generation of prompts that are closely aligned with a semantically similar prompts set ($P^{+}_{i}$) while being distinctly different from a semantically unrelated prompts set ($P^{-}_{i}$). The prompts sampling strategy is defined as follows: For each class $k$, global prompts for class $k$ are selected, and their cosine similarity with the locally generated prompt of $k$ is computed. For clients $\mathcal{U}_{o}$ and $\mathcal{U}_{n}$ (single domain), the closest prompt is the positive sample $P^{+}$. For clients $\mathcal{U}_{b}$ (two domains), the two closest prompts are positive samples, and all remaining prompts are negative samples $P^{-}_{i}$. Drawing inspiration from the FPL [15], where contrastive learning was used to augment the model’s proficiency in learning domain-invariant knowledge, we introduce domain-specific prompt contrastive learning with temperature decay specifically designed to enhance the adaptability of prompt generation in the CDAP. It is natural to derive the following optimization objective term:
$\mathcal{L}_{DPCL}=-\log\frac{\exp(\text{sim}(u_{i},P^{+}_{i})/\tau^{\prime})}{\exp(\text{sim}(u_{i},P^{+}_{i})/\tau^{\prime})+\exp(\text{sim}(u_{i},P^{-}_{i})/\tau^{\prime})}\text{.}$ (6)
Additionally, we address increasing domain diversity globally by proposing a method to calculate the contrastive temperature, $\tau$, which adaptively adjusts the loss. Initially allowing flexibility in distinguishing between positive and negative prompts, it becomes more stringent as learning progresses. The calculation of the updated temperature is as follows:
$\tau^{\prime}=\max\left(\tau_{\text{min}},\tau\cdot\left(1-\left(\gamma+(t-1)\cdot\beta\right)\right)\right)\text{,}$ (7)
where $\tau_{\text{min}}$ is the minimum temperature, $\gamma$ represents the base decay rate of the temperature, and $\beta$ is the rate of increment. Both $\gamma$ and $\beta$ are confined within the interval [0,1], indicating that their values range from 0 to 1, inclusive.
Local Domain-invariant Knowledge Learning with Prompting. To address the complexity of representing each class in our global clustering model due to varying cluster numbers, we average across all classes. This strategy allows us to create a generalized prompt $\bar{P}^{G}$ that effectively represents global domain information, ensuring a balanced and simplified approach to encapsulating the diversity of each domain, where
$\bar{P}^{g}=\left[AVG(\hat{P}^{g,1}),\dots,AVG(\hat{P}^{g,k})\right]\text{.}$ (8)
We enhance our model’s prediction robustness and mitigate forgetting by using global heterogeneous domain-specific prompts as diverse stimuli. This method helps the model learn domain-invariant knowledge. To support this, we introduce the $GPL$ loss, formulated by developing a CrossEntory [5] (CE) mechanism, which is defined by the following equation:
$\mathcal{L}_{GPL}=-\sum_{i=1}^{N}y_{i}\log\left(\text{softmax}\left(\xi_{g}\right)\right)\text{,}$ (9)
where $\mathcal{L}_{GPL}$ a loss function, sums over $N$ instances, using $y_{i}$ as the true label and applying softmax to the output of logits $\xi_{g}=G\left([\bar{P}^{g},h(x)]\right)$, classifier $G$ which combines the global prompt $\bar{P}^{g}$ with feature $h(x)$. Concurrently, locally CDAP generated prompts $P^{l}$ are integrated into the CE loss function, $\mathcal{L}_{CE}$, to capture local domain-specific features, where $\xi_{l}=G\left([P^{l},h(x)]\right)$, thus reinforcing the model’s discriminative capability within local domains. The formulation of this loss function is as follows:
$\mathcal{L}_{CE}=-\sum_{i=1}^{N}y_{i}\log\left(\text{softmax}\left(\xi_{l}\right)\right)\text{.}$ (10)
The final optimization objective is succinctly captured as:
$\mathcal{L}=\mathcal{L}_{CE}+\mathcal{L}_{GPL}+\mathcal{L}_{DPCL}\text{.}$ (11)
Our Algorithm 1 of the FDIL process is outlined in Appendix B.
4 Experiments
4.1 Experimental Setup
Datasets. We evaluate our framework on four widely used datasets for image classification tasks. The first, Digits-five [32, 33, 4], consists of 215,695 images across 10 classes and five domains, each image sized at 32x32 pixels. The second dataset, OfficeCaltech10 [35], contains 2,533 images distributed over 10 classes and four domains, also in 32x32 pixels. The third, PACS [19], includes 9,991 images across seven classes and four domains, each image being 224x224 pixels. Lastly, we use a subset of DomainNet [27], which we have named FedDomainNet, featuring 100,361 images in 48 classes across six domains, with a resolution of 224x224 pixels. All details of datasets can be found in Appendix C
Baselines. In our study, we benchmark RefFil against well-known rehearsal-free methods. Specifically, we compare it with the finetune approach, which involves straightforward model updates but is significantly impacted by catastrophic forgetting. Additionally, we consider four prominent rehearsal-free approaches: LwF [21], EWC [18], L2P [37] and DualPrompt [36]. These methods are fundamentally based on regularization methods.
However, they were originally proposed for centralized continual learning contexts, and consider to FDIL is an under-explored field. We adapted these methodologies for our FDIL setting, resulting in variants we refer to as FedLwF, FedEWC, FedL2P and FedDualPrompt, respectively. Particularly, to ensure a fair comparison, we initially deactivated the prompt pool feature in FedL2P and FedDualPrompt. Moreover, to demonstrate the superior efficacy of our RefFiL approach, even when competing methods employ prompt pool as a rehearsal technique, we also evaluate versions of FedL2p and FedDualPrompt with the prompt pool feature reactivated. These special versions are indicated with a superscript in the Table 1 3 2 4.
Evaluation Metrics. Our evaluation primarily includes top-1 accuracy, reflecting the model’s most probable correct prediction. We also report Average (Avg %) accuracy, as defined in iCaRL [30], which is the mean accuracy across all learning steps, indicating the model’s stability and consistency. Additionally, we present the Final (Last %) accuracy, measured after the last learning step, to assess the model’s knowledge retention and consolidation capabilities.
Digits-five OfficeCaltech10 PACS FedDomainNet
Methods Avg $\Delta$ Last $\Delta$ Avg $\Delta$ Last $\Delta$ Avg $\Delta$ Last $\Delta$ Avg $\Delta$ Last $\Delta$
Finetune 77.39 9.55 49.80 12.31 44.56 9.00 19.29 14.37 40.18 15.14 30.82 13.45 28.46 0.47 18.07 0.82
FedLwF 77.58 9.36 56.86 5.25 46.78 6.78 28.74 4.92 40.12 15.20 26.61 17.66 27.95 0.98 17.96 1.02
FedEWC 78.20 8.74 45.89 16.22 44.38 9.18 15.55 18.11 40.27 15.05 27.36 16.91 26.10 2.83 18.37 0.58
FedL2P 83.45 3.49 57.65 4.46 46.51 7.05 26.57 7.09 49.68 5.64 35.32 8.95 25.26 3.67 18.42 0.56
FedL2P 84.86 2.08 60.17 1.94 45.41 8.15 25.20 8.46 50.00 5.32 34.52 9.75 22.18 6.75 15.59 3.39
FedDualPrompt 85.15 1.79 59.30 2.81 45.15 8.41 23.82 9.84 54.05 1.27 41.07 3.20 28.25 0.68 18.05 0.93
FedDualPrompt 84.39 2.55 58.34 3.77 47.86 5.70 27.76 5.90 52.79 2.53 37.62 6.65 28.53 0.40 17.76 1.22
RefFiL 86.94 - 62.11 - 53.56 - 33.66 - 55.32 - 44.27 - 28.93 - 18.98 -
Table 1: Summarised results on four datasets: Digits-five with five incremental tasks, FedDomainNet with six, and both OfficeCaltech10 and PACS with four tasks. Displaying average accuracy (Avg%) and final task accuracy (Last%). The amount of change in our method compared to other methods is marked as $\Delta$, The results of our RefFiL are highlighted in colour
Means this baseline enabled prompt pool
Digits-five OfficeCaltech10 PACS FedDomainNet
Methods Avg $\Delta$ Last $\Delta$ Avg $\Delta$ Last $\Delta$ Avg $\Delta$ Last $\Delta$ Avg $\Delta$ Last $\Delta$
Finetune 59.84 9.52 58.20 2.64 37.60 6.73 25.20 13.19 46.99 4.09 38.97 7.75 31.85 1.49 11.58 4.16
FedLwF 65.22 4.14 59.36 1.48 38.76 5.57 25.20 13.19 43.43 7.65 30.17 16.55 31.33 2.01 11.01 4.73
FedEWC 64.00 5.36 59.54 1.30 38.26 6.07 27.95 10.44 43.60 7.48 30.22 16.50 30.38 2.96 12.03 3.71
FedL2P 66.00 3.36 59.84 1.00 41.58 2.75 34.45 3.94 45.99 5.09 31.02 15.70 25.19 8.15 9.51 6.23
FedL2P 64.45 4.91 59.74 1.10 41.24 3.13 31.50 6.89 45.39 5.69 35.42 11.30 22.95 10.39 7.32 8.42
FedDualPrompt 65.31 4.05 60.94 0.10 40.47 3.86 31.50 6.89 48.41 2.67 42.32 4.40 33.09 0.25 14.54 1.20
FedDualPrompt 66.61 2.75 60.94 0.10 39.73 4.6 30.91 7.48 47.64 3.44 42.82 3.90 30.11 3.23 14.54 1.20
RefFiL 69.36 - 60.84 - 44.33 - 38.39 - 51.08 - 46.72 - 33.34 - 15.74 -
Table 2: Summarised results in new domain order of Table 1.
Task 1 $\rightarrow$ 5 on Digit-Five Task 1 $\rightarrow$ 4 on OfficeCaltech10
Methods MNIST MNIST-M USPS SVHN SYN Avg Amazon Caltech Webcam DSLR Avg
Finetune 99.68 97.75 63.87 75.84 49.80 77.39 76.56 57.79 24.58 19.29 44.56
FedLwF 99.68 92.80 69.16 69.39 56.86 77.58 76.56 53.24 28.57 28.74 46.78
FedEWC 99.68 97.48 74.63 73.32 45.89 78.20 76.56 56.59 29.83 15.55 44.38
FedL2P 99.66 98.06 80.01 81.89 57.65 83.45 76.56 51.80 31.09 26.57 46.51
FedL2P 99.64 97.65 85.18 81.65 60.17 84.86 71.35 55.88 29.20 25.20 45.41
FedDualPrompt 99.67 97.96 86.88 81.95 59.30 85.15 74.48 50.36 31.93 23.82 45.15
FedDualPrompt 99.65 97.90 84.68 81.40 58.34 84.39 75.90 53.96 33.82 27.76 47.86
RefFiL 99.68 98.25 90.96 83.70 62.11 86.94 78.65 61.15 40.76 33.66 53.56
Task 1 $\rightarrow$ 6 on FedDomainNet Task 1 $\rightarrow$ 4 on PACS
Methods Clipart Infograph Painting Quickdraw Real Sketch Avg Photo Cartoon Sketch Art Painting Avg
Finetune 51.48 15.89 28.05 27.84 29.45 18.07 28.46 61.68 47.45 36.12 30.82 40.18
FedLwF 51.48 18.10 26.71 25.98 27.47 17.96 27.95 61.68 47.07 25.11 26.61 40.12
FedEWC 50.76 15.46 22.66 21.87 27.45 18.37 26.10 63.17 47.70 23.66 27.36 40.27
FedL2P 40.55 13.19 21.09 28.15 30.13 18.42 25.26 64.97 48.32 50.09 35.32 49.68
FedL2P 37.63 9.29 16.79 27.09 26.68 15.59 22.18 65.57 54.67 45.25 34.52 50.00
FedDualPrompt 51.17 19.48 28.74 22.68 29.40 18.05 28.25 73.65 56.54 44.93 41.07 54.05
FedDualPrompt 51.14 20.20 28.91 23.09 30.07 17.76 28.53 75.75 54.55 43.23 37.62 52.79
RefFiL 51.27 20.91 29.23 22.57 30.62 18.98 28.93 73.95 59.90 43.17 44.27 55.32
Table 3: Comparison of RefFiL’s performance with five baseline methods on four widely used datasets, showcasing average accuracy (Avg %) and accuracy for each domain task (%). At each task step, all clients have an equal number of classes but quantity shift data.
Task 1 $\rightarrow$ 5 on Digit-Five Task 1 $\rightarrow$ 4 on OfficeCaltech10
Methods SVHN MNIST SYN USPS MNIST-M Avg Caltech Amazon DSLR Webcam Avg
Finetune 94.97 58.35 49.04 38.66 58.20 59.84 49.78 58.27 17.15 25.20 37.60
FedLwF 94.97 73.21 54.73 43.82 59.36 65.22 49.78 57.79 22.27 25.20 38.76
FedEWC 95.03 64.32 50.22 50.88 59.54 64.00 48.00 56.83 20.27 27.95 38.26
FedL2P 94.85 73.54 53.19 48.56 59.84 66.00 49.78 58.03 24.05 34.45 41.58
FedL2P 94.80 73.45 51.07 43.21 59.74 64.45 50.67 58.27 24.50 31.50 41.24
FedDualPrompt 94.78 70.71 54.06 46.04 60.94 65.31 48.00 58.75 23.61 31.50 40.47
FedDualPrompt 94.65 77.02 54.43 46.01 60.94 66.61 50.22 57.07 20.71 30.91 39.73
RefFiL 95.35 76.03 59.90 54.68 60.84 69.36 52.00 63.31 23.61 38.39 44.33
Task 1 $\rightarrow$ 6 on FedDomainNet Task 1 $\rightarrow$ 4 on PACS
Methods Infograph Sketch Quickdraw Real Painting Clipart Avg Cartoon Photo Sketch Art Painting Avg
Finetune 68.84 33.94 28.94 26.12 21.73 11.58 31.85 68.23 40.97 39.77 38.97 46.99
FedLwF 68.84 34.87 28.82 23.88 20.53 11.01 31.33 68.23 36.11 39.21 30.17 43.43
FedEWC 68.11 34.66 24.63 24.10 18.75 12.03 30.38 69.94 38.23 36.00 30.22 43.60
FedL2P 53.39 26.76 27.57 17.92 15.98 9.51 25.19 68.23 42.34 42.73 31.02 45.99
FedL2P 51.89 24.86 26.37 14.64 12.62 7.32 22.95 66.95 44.71 34.49 35.42 45.39
FedDualPrompt 70.67 41.50 25.82 25.30 20.73 14.54 33.09 69.94 41.34 40.03 42.32 48.41
FedDualPrompt 70.33 40.63 25.69 25.48 21.40 14.54 30.11 66.74 41.72 39.27 42.82 47.64
RefFiL 69.02 42.48 24.70 26.19 21.93 15.74 33.34 73.13 45.33 39.14 46.72 51.08
Table 4: Comparison of RefFiL’s performance in new domain order of table 3.
Implementation details. In our study, all methods are implemented using PyTorch [26] and run on a single NVIDIA RTX 4090 GPU, using ResNet10 [14] as the feature extractor classification model backbone. We designed a simple embedding model as the feature map tokenizer, similar to ViT [10], with initialized-only and frozen parameters for feature embedding. The FedAvg [24] algorithm was employed for global model aggregation. Each task involves training the model over $R=30$ rounds. The number of epochs for local updates on each client is set at 20. We use learning 0.04 for FedDomainNet, 0.06 for OfficeCaltech10 and 0.03 for the rest of the datasets to get the proper performance. Additionally, we use SGD as an optimizer for all experiments. Unless specified otherwise, the constraint factor $\lambda$ in the EWC method is fixed at 300, and the temperature for distillation is set to the default value of 2. Our client increment setting aligns with what we outlined in Appendix A. Specifically, for experiments on the Digit-Five, PACS, and FedDomainNet datasets, we start with 20 clients, select 10, and increment by 2 clients with each new task. Conversely, for experiments on the OfficeCaltech10 dataset, considering the data sample limitation, we start with 10 clients, select 5, and increment by 1 client per new task. In all experiments, 80% of the $\mathcal{U}o$ old clients transition to a new task for each respective task. For the hyperparameters in Eq. 7, we fix $\tau_{\text{min}}$ at 0.3, $\gamma$ at 0.1, $\beta$ at 0.05, and $\tau$ at 0.9.
4.2 Qualitative Results
In our study, we evaluated the performance of seven baseline methods alongside our proposed method, RefFiL, across four widely used datasets: Digit-Five, OfficeCaltech10, FedDomainNet, and PACS. The results, detailed in Table 1, Table 3, Table 2, and Table 4, indicate that RefFiL consistently outperforms the baseline methods in both averaged and last task accuracy, even under the impact of catastrophic forgetting. Notably, RefFiL maintains its advanced performance consistently, regardless of changes in domain order. To provide a comprehensive comparison, we conducted three additional experiments, which are discussed in Appendix D. These experiments illustrate the accuracy distribution across different domains as tasks change and analyze the decision-making boundaries of all baseline methods.
RefFiL efficiently mitigates forgetting. We report detailed performance in Table 3. The means the baseline enabled prompt pool to save generated prompts and then be sampled one for better mitigate forgetting. And the result can be summarised in Table 1, which demonstrates that RefFiL consistently achieves state-of-the-art (SOTA) results, showcasing robust performance across all four datasets in terms of average accuracy (Avg) and performance on the final task of each dataset (Last). Specifically, RefFiL shows a remarkable outperforming over the baseline methods in different scenarios. In smaller and less complex datasets like Digit-Five, RefFiL surpasses other baselines by 1.79% to 9.55% in average accuracy and by 1.94% to 16.22% in the final task accuracy. In the OfficeCaltech10 dataset, RefFiL outperforms baselines by 5.70% to 9.18% in average accuracy and 5.90% to 18.11% in the last task accuracy. Moreover, in larger and more complex image datasets like PACS, RefFiL continues to lead, achieving 1.27% to 15.20% higher average accuracy and 3.20% to 17.66% higher accuracy in the final task compared to the baselines. Similarly, despite the challenges posed by the FedDomainNet dataset, such as sparse data distribution across 48 classes and high-resolution images. RefFiL still demonstrated improvement. RefFiL outshines others with an improvement of 0.40% to 6.75% in average accuracy and 0.56% to 3.39% in the final task accuracy. These results collectively underscore the effectiveness and adaptability of RefFiL in handling diverse and challenging datasets, maintaining consistently high performance both overall and in the crucial final tasks.
RefFiL produces domain order-independent consistency results.
To demonstrate the consistency of our results across different domain introduction orders, we generated new results presented in Table 4 and summarized in Table 2. In these experiments, we randomly altered the order of domain introductions. Specifically, in Table 4, the domain orders for Digit-Five, OfficeCaltech10, and FedDomainNet were significantly disrupted, while in PACS, only the order of the first two domains was changed.
This setup highlights the robustness of our method, showing that RefFiL maintains high performance despite variations in the sequence of domain introductions. Compared to the results in Table 1, although FedDualPrompt and FedDualPrompt achieved better performance in the Last metric, RefFiL consistently outperformed these methods in the more critical Avg metric, indicating its superior overall performance. For both the OfficeCaltech10 and PACS datasets, RefFiL’s advantages in Avg and Last metrics were magnified compared to most of the baselines. In the most challenging dataset, FedDomainNet, while RefFiL’s advantage in Avg reduced from 0.40% to 0.20%, its Last metric slightly improved from 0.93% to 1.20%. These results significantly amplified RefFiL’s advantages over other baselines.
These results illustrate that RefFiL efficiently and stably mitigates catastrophic forgetting by enhancing the model’s ability to learn from domain-invariant knowledge. The consistent high performance across various domain orders underscores RefFiL’s robustness and adaptability in FDIL scenarios.
4.3 Ablation Study
Methods CDAP GPL DPCL Avg $\Delta$ Last $\Delta$
44.56 - 19.29 -
49.78 5.22 27.56 8.27
RefFiL 47.94 3.38 26.38 7.09
50.32 5.76 25.39 6.10
49.45 4.89 30.12 10.83
53.56 9.00 33.66 14.37
Table 5: Ablation study of RefFil on OfficeCaltech10 dataset, and the amount of improvement in RefFiL with different components compared to baseline (Finetune) is marked as $\Delta$.
To understand how different components can mitigate catastrophic forgetting problems, we conduct the experiments as shown in Table 5; we dissect the impact of each module within RefFil, as well as their synergistic effects. Each experimental configuration involving a separate component is indicated by a checkmark (✓).
Our findings reveal that using either the CDAP or GPL module individually enhances both average and final task accuracy, but their combination results in only a slight improvement in average accuracy compared to when they are used separately. Considering that the DPCL module cannot function in isolation, we explored its effectiveness in conjunction with the GPL module. This combination markedly improved the model’s performance, especially in the accuracy of the final task. Most importantly, when all three core components CDAP, GPL, and DPCL are integrated, we observe a substantial improvement in both average and last-task accuracy.
This comprehensive ablation study confirms the essentiality and complementarity of each module. Their collective operation significantly enhances the global model’s performance in domain-incremental scenarios. The results validate the necessity of each component, underscoring their collective contribution to achieving optimal outcomes in FDIL.
5 Conclusion
In this paper, we introduced RefFil, a rehearsal-free federated domain-incremental learning framework designed to augment cross-domain prediction robustness by enhancing the acquisition of domain-invariant knowledge. Our approach includes a client-wise domain adaptive prompt generator for creating personalized, instance-level prompts, a global prompt learning paradigm for breaking down domain-specific information silos, and a novel domain-specific prompt contrastive learning method with temperature decay for discriminating between relevant and irrelevant prompts. Through extensive experiments, we show that RefFil framework excels in capturing domain invariant knowledge mitigating catastrophic forgetting in federated domain-incremental learning. It is also important to acknowledge that while RefFil is proposed for learning domain-invariant knowledge through prompting, it may also contribute to the development of fine-tuning large pretrained models in domain-specific areas.
Appendix A Preliminaries
Federated Domain-incremental Learning. In standard domain-incremental learning [25], a series of sequential tasks are denoted as $\mathcal{T}=\{\mathcal{T}^{t}\}_{t=1}^{T}$, with $T$ indicating the total number of tasks. Each task $\mathcal{T}^{t}$, representing the $t$-th task in the sequence, consists of $N^{t}$ data-label pairs, represented as $\mathcal{T}^{t}=\left\{x_{i}^{t},y_{i}^{t}\right\}_{i=1}^{N^{t}}$. Although each task is associated with a distinct domain $\mathcal{D}^{t}$, all tasks share the same label space $\mathcal{Y}$. The entire domain space across all tasks is represented as $\mathcal{D}=\{\mathcal{D}^{t}\}_{t=1}^{T}$, with the data distribution being completely inconsistent between different task domains, indicated by $\mathcal{D}^{t}\neq\mathcal{D}^{t-1}$.
We adapt traditional domain-incremental learning in the context of FDIL. In this setup, there are $M$ local participants, denoted as =$\mathcal{M}=\left\{\mathcal{M}_{m}\right\}_{m=1}^{M}$, and a central global server $\mathcal{M}_{G}$. During each global communication round $r$ (where $r=1,\dots,R$), a subset of these local participants is randomly chosen for gradient aggregation. When a local participant $\mathcal{M}_{l}$ is selected for a given global round in the $t$-th incremental task, it receives the latest global model $\theta^{{r},t}$.
The participant then trains this latest model on its local dataset $\mathbf{D}^{t}\sim\mathcal{P}_{l}^{\left|\mathbf{D}^{t}\right|}$ specific to the $t$-th task, where $\mathcal{T}_{m}^{t}=\left\{x_{mi}^{t},y_{mi}^{t}\right\}_{i=1}^{N_{l}^{t}}\subset\mathcal{T}^{t}$ consists of data from new domains, and $\mathbf{P}_{m}$ is the data quantity distribution of the $m$-th participant. These local datasets $\left\{\mathbf{P}_{m}\right\}_{m=1}^{M}$ are not independent and identically distributed (non-iid), showcasing a type of quantity shift in our setting. Each incremental task contains $K^{t}$ classes, maintaining consistency in class numbers across tasks ($K^{1}=K^{2}=\dots=K^{t}$) and labels ($\mathbf{y}^{1}=\mathbf{y}^{2}=\dots=\mathbf{y}^{t}$), but with varying domain and data distributions ($\mathbf{x}^{1}\neq\mathbf{x}^{2}\neq\dots\neq\mathbf{x}^{t}$) for each task $t\in\mathcal{T}$. Once the local participant $\mathcal{M}_{m}$ loads $\theta^{r,t}$ and completes training on the $t$-th task, it obtains an updated local model, denoted as $\theta^{r,t}_{l}$. All selected participants then transmit their local updated model to the central global server $\mathcal{M}_{G}$. Here, these models through aggregation form the updated global model $\theta^{(r+1),t}$ for the next round. Subsequently, the global server $\mathcal{M}_{G}$ broadcasts the parameters of $\theta^{r,t}$ back to all participants, marking the beginning of the next global communication round.
Client increment strategy. We divide local participants into three dynamic groups across each incremental task, comprising Old $\mathcal{U}_{o}$, In-between $\mathcal{U}_{b}$, and New $\mathcal{U}_{n}$. Group $\mathcal{U}_{o}$, with $M_{o}$ members, is restricted to working solely on data from previous domains, receiving no new domain data for the current task. Group $\mathcal{U}_{b}$, consisting of $M_{b}$ participants, handles both new domain data from the current task and data from previous tasks. The group $\mathcal{U}_{n}$, comprising $M_{n}$ new participants, focuses exclusively on the new domain data of the current task. These groups adapt dynamically with the arrival of incremental tasks. Specifically, the composition of $\mathcal{U}_{o},\mathcal{U}_{b},\mathcal{U}_{n}$ is randomly redefined at each global round, with $\mathcal{U}_{n}$ being irregularly supplemented at any round in the FDIL process. This approach gradually increases the total number of participants, $M=M_{o}+M_{b}+M_{n}$, as streaming tasks progress.
Learning with Prompts. Prompt techniques, initially developed for task adaptation in natural language processing [22], have predominantly been applied to transformer-based large models. However, these techniques can also be adapted for models using convolutional neural network (CNN) based feature extractors. This adaptation is particularly relevant in federated learning environments, where participant devices are more likely to be resource-constrained.
In scenarios where a CNN-based model serves as the backbone for prompt-based learning, the model architecture typically includes a CNN-based feature extractor $h$, patch embedding layer, attention block $b$, and the classifier $G$. Specifically, feature map $F$ is extracted by a feature extractor from an input image $x$ as $F=h(x)$. Let $F$ be split into $n$ $d$-dimensional patch tokens $\mathrm{PT}$ through the patch embedding layer. Then, a $d$-dimensional trainable class token [CLS] is attached at the beginning of patch tokens along the sequence length dimension, thereby creating the initial set of input tokens $\mathbf{I}\in\mathbb{R}^{(n+1)\times d}$ as
$\mathbf{I}=\left[\mathrm{CLS};\mathrm{PT}_{1},\dots,\mathrm{PT}_{n}\right]\in\mathbb{R}^{(n+1)\times d}.$ (12)
Following this, $\mathbf{I}$ is forwarded to the $b$-th attention block, which generates the input $\mathbf{I}_{b+1}$ for the subsequent attention block ($b+1$) as:
$\mathbf{I}_{b+1}=\operatorname{LN}\left(\mathbf{I}_{b}^{\prime}+\mathbf{I}_{b}^{\prime\prime}\right),$ (13)
where LN is layer normalization, $\mathbf{I}_{b}^{\prime\prime}=\operatorname{MLP}\left(\mathbf{I}_{b}^{\prime}\right)$ is multilayer perception (MLP), and $\mathbf{I}_{b}^{\prime}=\operatorname{LN}\left(\operatorname{MHSA}\left(\mathbf{I}_{b},\mathbf{I}_{b},\mathbf{I}_{b}\right)\right)$ is multi-head self-attention (MHSA) mechanism. This indicates that each attention block involves an MHSA mechanism, succeeded by an MLP with the skip connection [14], and an LN [1] (see Fig. 2). The classifier, which is a single feed-forward layer, determines the class labels. It does this by mapping the class token [CLS] from the output of the final attention block as:
$\hat{y}=G([CLS]_{B}),$ (14)
where $\hat{y}$ is the predicted distribution of class probabilities.
Appendix B Algorithm
Input : Total tasks $T$, Global communication round $R$, client local epoch $E$, number of participants $M$, participant local data $\mathbf{D}_{m}(x,y)$, participant local model $\theta_{m}$, learning rate $\eta$
Output : Trained global model $\theta^{R}$
1
2Central Server
3       Initialise global model $\theta^{r}$
4      for $t=1$ to $T$ do
5             for $r=1$ to $R$ do
6                   Participants Random Selection
7                  Broadcast $\theta^{r,t},\hat{P}^{g}$rightarrow Selected Participants
8                  $\theta^{r}_{m}$, $LPG^{k}_{m}\leftarrow$ Selected Participants
9                  $\theta^{r+1}\leftarrow\sum_{m=1}^{N}\frac{|\mathbf{D}^{t}_{m}|}{|\mathbf{D}^{t}|}\theta_{m}^{r}$
10                  $\hat{P}^{g,k}=\left\{LPG_{m}\right\}_{m=1}^{C}\xrightarrow{\Psi}\left\{LPG_{m}\right\}_{m=1}^{N}$ Eq.5 $\triangleright$ Global prompt clustering
11                  $\hat{P}^{g}=\left[\hat{P}^{g,1},\dots,\hat{P}^{g,k}\right]$
12
13            $\mathcal{U}_{n},\mathcal{U}_{b},\mathcal{U}_{o}\leftarrow$ $M$Sampling $\triangleright$ Participant task transfer sampling and increment
14            for $m$ in $\mathcal{U}_{b}$ do
15                   $\mathbf{D}^{t}_{m}=\text{concat}(\mathbf{D}^{t-1}_{m},\mathbf{D}^{t}_{m})$
16
17
18
19Participants
20       $P_{m}=\left\{\right\}$triangleright Initialise local prompts collection
21      $\theta^{r,t},\hat{P}^{g}$leftarrow Central Server
[…]
23      for $e=0,...,E$ do
24             for $x_{i},y_{i}\in\mathbf{D}^{t}_{m}$ do
25                   $u_{i},p_{l},\xi_{l},\xi_{g}=\theta^{r,t}(x_{i},\bar{P}^{g})$
26                  $\mathcal{L}_{CE}\leftarrow$ Eq.10 $(u_{i},\xi_{l})$
27                  $\mathcal{L}_{GPL}\leftarrow$ Eq.9 $(u_{i},\xi_{g})$
28                  $\mathcal{L}_{DPCL}\leftarrow$ Eq.6 $(p_{l},\hat{P}^{g})$
29                  $\mathcal{L}=\mathcal{L}_{CE}+\mathcal{L}_{GPL}+\mathcal{L}_{DPCL}$
30                  $\theta^{r}_{m}\leftarrow\theta^{r}_{m}-\eta\nabla\mathcal{L}(\theta^{r}_{m};b)$
31                  if $e=(E-1)$ then
32                         $P_{m}[k]$ append $p_{i}$
33
34
35
36      for $k$ in $P_{m}$ keys do
37             $\triangleright$ Average local prompts into a representative prompt $LPG_{m}=\left[\frac{1}{N}\sum_{j=1}^{N}P^{1}_{m,j},\dots,\frac{1}{N}\sum_{j=1}^{N}P^{k}_{m,j}\right]$
38
39      Send $\theta^{r}_{m}$, $LPG_{m}$rightarrow Central Server
Algorithm 1 Rehearsal-free federated domain-incremental learning
Appendix C Datasets
We evaluate our framework on four widely used datasets for image classification tasks, Digits-five [32, 33, 4], OfficeCaltech10 [35], PACS [19], and a subset of DomainNet [27], we named FedDomainNet.
  • Digits-five [32, 33, 4]: This dataset is a comprehensive collection of five most popular digit datasets for digit recognition task, encompassing 10 categories (digits 0-9). It aggregates five distinct domains: MNIST (55000 samples), MNIST-M (55000 samples), Synthetic Digits (25000 samples), SVHN (73257 samples), and USPS (7438 samples). In total, it comprises 215,695 images, each of 32x32 pixels.
  • OfficeCaltech10 [35]: This dataset spans four domains: Amazon (958 samples), Caltech(1123 samples), Webcam(295 samples), and DSLR(157 samples). It features 10 classes across each domain, culminating in a total of 2533 images, resized to 32x32 pixels for uniformity.
  • PACS [19]: Encompassing four domains — Photo (1,670 images), Art Painting (2,048 images), Cartoon (2,344 images) and Sketch (3,929 images) — this dataset is distinguished by its varied artistic representations. Each domain includes 7 categories, amounting to a total of 9,991 images. Originally of size 227x227, these images are cropped to 224x224 pixels to standardize the dataset.
  • FedDomainNet: This dataset, a curated subset derived from DomainNet [27], features six domains: Clipart (9749 samples), Infograph (11010 samples), Painting (15228 samples), Quickdraw (24000 samples), Real (26476 samples), and Sketch (13898 samples). Each domain offers 48 classes, resulting in a comprehensive collection of 100,361 images. For consistency in analysis, these images are also cropped to a size of 224x224 pixels.
C.1 FedDomainNet Dataset Details
In terms of dataset processing, we address the challenge posed by the vast scale of the DomainNet database [27], which comprises an extensive collection of 24 divisions with 345 categories, totaling 586,575 images. Such a large dataset presents considerable computational demands in the context of federated learning. Specifically, for federated domain-incremental learning, the key objective is to ensure each client sustains a diverse range of categories while possessing sufficient domain variety to mimic a continual data stream. To meet this requirement, we strategically extracted a representative subset from the cleaned version of the original DomainNet dataset for our experiments. As outlined in Table 6, from each of the 24 divisions, we carefully selected two classes. These classes were chosen based on their balanced representation across six domains and their clear categorical distinctions, resulting in a total of 48 classes. The subset from the original version comprised 102,986 images; however, utilizing the cleaned version, our final dataset consisted of 100,361 images. This selection was aimed at maintaining a uniform global distribution Figure 3 of types and data volumes to underscore the influence of domain shifts on the model.
Figure 3: FedDomainNet data distribution statistical information
Division Class clp inf pnt qdr rel skt Total
Furniture teapot 222 209 391 500 631 327 2280
streetlight 326 113 537 500 463 268 2207
Mammal tiger 315 285 422 500 607 386 2515
whale 343 432 357 500 671 272 2575
Tool stethoscope 343 107 346 500 496 237 2029
sword 139 124 470 500 591 384 2208
Cloth shoe 127 291 260 500 587 645 2410
bracelet 293 123 150 500 715 300 2081
Electricity headphones 285 224 181 500 551 188 1929
toaster 196 337 107 500 536 267 1943
Building golf club 207 169 650 500 552 695 2773
windmill 245 372 397 500 635 245 2394
Office cup 128 52 582 500 406 396 2064
map 42 206 423 500 507 193 1871
Human Body goatee 255 236 129 500 562 219 1901
eye 108 168 292 500 695 489 2252
Road Transportation train 109 373 406 500 681 240 2309
tractor 154 316 183 500 636 263 2052
Food bread 197 232 315 500 794 276 2314
ice cream 160 187 311 500 657 184 1999
Nature sun 248 352 572 500 161 258 2091
tornado 169 329 373 500 497 211 2079
Cold Blooded sea turtle 236 190 410 500 621 254 2211
fish 130 195 429 500 479 373 2106
Music guitar 103 204 203 500 632 183 1825
trombone 227 195 175 500 484 191 1772
Fruit strawberry 357 308 530 500 454 198 2347
watermelon 193 401 410 500 671 128 2303
Sport snorkel 278 81 179 500 689 397 2124
yoga 165 447 161 500 371 251 1895
Tree tree 126 511 571 500 536 555 2799
flower 253 140 485 500 360 336 2074
Bird bird 336 208 222 500 803 306 2375
penguin 121 201 447 500 700 209 2178
Vegetable mushroom 136 298 254 500 788 252 2228
broccoli 105 229 100 500 679 181 1794
Shape zigzag 323 412 110 500 515 144 2004
triangle 183 364 298 500 376 303 2024
Kitchen spoon 228 127 158 500 534 406 1953
hourglass 100 100 206 500 289 134 1329
Water Transportation sailboat 162 119 322 500 422 361 1886
submarine 344 183 550 500 607 207 2391
Sky Transportation helicopter 145 216 257 500 804 200 2122
hot air balloon. 198 48 453 500 732 170 2101
Insect bee 202 233 313 500 452 144 1844
butterfly 160 162 387 500 658 249 2116
Others feather 268 432 344 500 505 336 2385
snowman 174 123 901 500 114 712 2524
Total - 9864 11364 16729 24000 26906 14123 102986
Total (Cleaned) - 9749 11010 15228 24000 26476 13898 100361
Table 6: Detailed statistics of the FedDomainNet dataset
Appendix D Additional Experiments and Results
RefFiL performance stable accuracy distribution across different domains. We also plot the accuracy distribution for each domain of all tasks, presented in Figure 4, illustrating the distribution of accuracy across various baseline methods for different tasks and domains. The subplots show a clear comparison between the baseline methods, with each box plot providing insights into the median accuracy, interquartile range, and variability for each method.
The subplot corresponding to our ReFiL stands out due to its narrow interquartile range and high median accuracy across all tasks, indicating a consistent and reliable performance. For example, in the MNIST task, ReFiL achieves a median accuracy of 99.64%, with minimal variability among the accuracy values. This consistency is also observed in more complex tasks like MNIST-M and USPS, where ReFiL maintains a high median accuracy compared to other baselines.
Additionally, ReFiL’s box plot shows fewer outliers than other baselines, indicating that its performance is less susceptible to variability or domain shift. This resilience is especially notable in challenging tasks like SVHN, where ReFiL still achieves a median accuracy of 95.18%, outperforming several other baselines. Overall, ReFiL’s superior performance in terms of both high accuracy and low variability is consistently demonstrated across all tasks, highlighting its robustness and effectiveness as a baseline method.
RefFiL clearer decision-making boundaries. Additionally, to demonstrate the model’s capability to learn a generalizable decision boundary across various class domains, we present a t-SNE visualization in Figure 5, comparing five baselines with our RefFiL on the Digits-Five Dataset. This visualization is particularly effective in showing how well the model differentiates between class domains at each task stage. The incremental task structure is as follows: Task 1 evaluates the model using only the MNIST dataset. Task 2 tests the model on a combined dataset of MNIST and MNIST-M. Task 3 expands the assessment to include MNIST, MNIST-M, and USPS. In Task 4, the model faces a dataset that integrates MNIST, MNIST-M, USPS, and SYN. Finally, Task 5 challenges the model with all previously learned datasets, adding SVHN into the mix. Compared to other baseline methods, RefFiL demonstrates a superior performance starting from Task 2, as evidenced by the greater clarity and distinctness of each cluster’s boundaries in the t-SNE visualization. Notably, the key areas marked by red boxes, which indicate regions where different classes overlap or are challenging to separate, are consistently smaller with RefFiL. This indicates that RefFiL is more effective at distinguishing between classes, leading to less confusion and more precise clustering than the baselines.
More specifically, in Figure 5, we present a detailed t-SNE visualization of Task 5, illustrating the global model’s performance after the completion of this task. This figure evaluates the task 5 output model’s performance in handling all learned domain datasets ranging from MNIST to SVHN. It is evident from the visualization that RefFiL achieves a more distinct decision boundary, particularly for the MNIST, USPS, and SYN datasets. Moreover, for the MNIST-M and SVHN datasets, RefFiL demonstrates significant diffusion of cluster boundaries. This diffusion is indicative of RefFiL’s proficiency in achieving effective class separation across a diverse range of domain datasets.
Figure 4: Box plots of the accuracy distribution for each domain in every task.
Figure 5: t-SNE Visualization of Five Baselines and Our RefFiL Model Across Task Steps in the Digits-Five Dataset: This illustration captures the global model’s ability in learning a generalizable decision boundary at each task step, evaluated across all previously learned domains. The greater the clarity and diffusion of each cluster’s boundaries, the better. The core areas where different classes are challenging to distinguish are highlighted with red boxes. Smaller box sizes indicate fewer data points that are difficult to cluster.
Figure 6: Detailed t-SNE Visualization of Task 5 in Figure5 for Five Baselines and Our RefFiL Model. This figure showcases the global model, updated post Task 5, evaluated separately on datasets from MNIST to SVHN, reflecting the learning progression from Task 1 to 5. Enhanced clarity and diffusion of each cluster’s boundaries signify the global model’s proficiency in developing generalizable decision boundaries. For detailed description, refer to Appendix D.
Appendix E Limitations
While our method significantly mitigates forgetting problems in a rehearsal-free manner, qualitative results reveal a notable limitation when dealing with the more complex FedDomainNet dataset. Additionally, our approach relies on the prompt encoder leveraging the client’s local task ID as a reference. In situations where the task ID is not uniquely managed globally and the client is unaware of it, this dependency can become a limitation. Furthermore, although RefFiL is an efficient method for learning domain-incremental data, federated learning from streaming data presents the additional challenge of sequentially learning from both new domains and new classes. Addressing these challenges will be crucial for enhancing the scalability and applicability of RefFiL in more dynamic and complex federated learning environments.
Appendix F Abbreviations List
Symbol Description Symbol Description
$T$ Tasks $\mathcal{T}$ Set of tasks
$t$ Current task $x$ Input data (image)
$y$ Input label $\mathcal{Y}$ Label space
$\mathcal{D}$ Domains space $\mathcal{D}^{t-1}$ Previous task domain
$U$ Individual participant $U_{G}$ Central server
$R$ Global communication round $r$ Single global communication round
$K$ Total numbers of class $\mathbf{P}$ Data quantity distribution
$\mathcal{U}$ Participants group $\mathcal{U_{n}}$ New task participants group
[…]
$h$ Feature Extractor $F$ Feature Map
$G$ Classifier $b$ Attention Block
$B$ Set of attention blocks $\mathrm{PT}$ Patch Token
$\mathbf{I}$ Set of input tokens $\operatorname{MHSA}$ Multi-head self-attention
$\operatorname{MLP}$ Multi-layer perception $\operatorname{CCDA}$ Cross-client domain adaptation
$\operatorname{LN}$ Layer normalization $\operatorname{LT}$ Feature-Wise Linear Modulation
Transformation Layer
$\mathbf{D}$ Dataset $\mathcal{G}$ CDAP generator
$\phi$ Linear layer $\alpha$ Affine transformation parameter
$\lambda$ Affine transformation parameter $v$ Conditional input embedding
$P$ Prompts $P^{g}$ Plobal prompts
$P^{m}$ Local prompts $\hat{P}^{g}$ Global clustered prompts
$\bar{P}^{g}$ Global averaged clustered prompts $\tau$ Base temperature
$\gamma$ Base temperature decay rate $\beta$ Decay rate increment rate
$\tau^{\prime}$ Decayed temperature $P^{+}$ Positive prompts
$P^{-}$ Negative prompts $E$ client local epoch
$M$ Number of participants $\eta$ Learning rate
$\Psi$ Cluster model $\xi_{l}$ Logits local prompt as input
$\xi_{g}$ Logits global prompt as input $\mathcal{L}_{CE}$ Classification cross-entropy loss
function
$\mathcal{L}_{GPL}$ Global prompt learning loss function $\mathcal{L}_{DPCL}$ Domain-specific prompt contrastive
learning loss function
$\mathcal{L}$ Main loss function $CLS$ ViT class token
CDAP Client-wise Domain Adaptive Prompt GPL Global Prompt Learning loss
DPCL Domain-specific Prompt Contrastive Learning loss LPG Local Prompts Group